システム基礎講座 図書目録

洋書

平成 7 年 10 月 18 日

A [AB93] M. L. Abel, J. P. Braselton, Differential Equations with Mathematica, AP Professional, 1993. [0-12-041539-9] 94 年 5 月購入

[AMO93] R. K. Ahuja, T. L. Magnanti, J. B. Orlin, Network Flows: Theory, Algorithms, and Applications, Prentice-Hall, 1993. [0-13-027765-7] 93 年 7 月購入

[Aka94] T. J. Akai, Applied Numerical Methods for Engineers, John Wiley & Sons, 1994. [0-471-02017-6] 94 年 11 月購入

[AHT92] M. Akgül, H. W. Hamacher, S. Tüfekçi, Combinatorial Optimization: New Frontiers in Theory and Practice, Springer, 1992. 93 年 6 月購入

[AG90] E. L. Allgowerm, K. Georg, Numerical Continuation Methods: An Introduction, Springer-Verlag, 1990. [3-540-12760-7] 94 年 8 月購入

[Arb93] A. Arbel, Exploring Interior-Point Linear Programming: Algorithms and Software, The MIT Press, 1993. [0-262-51073-1] 94 年 2 月購入

[Aub93] J-P. Aubin, Optima and Equilibria: An Introduction to Nonlinear Analysis, Springer-Verlag, 1993. [3-540-52121-6] 94 年 2 月購入

B [Baa88] S. Baase, Computer Algorithms: Introduction to Design and Analysis, second edition, Addison-Wesley, 1988. [0-201-06035-3] 94 年 7 月購入 [BC95] E. Balas, J. Clausen (eds.),

Integer Programmming and Combinatorial Optimization, Springer-Verlag, 1995. [3-540-59408-6] 95 年 7 月購入

[BC95] E. Balas, J. Clausen (eds.), Integer Programmming and Combinatorial Optimization, Springer-Verlag, 1995. [3-540-59408-6] 95 年 9 月購入 (2 冊目)

[BMMN95] M. O. Ball, T. L. Magnanti, C. L. Monma, G. L. Nemhauser (eds.), Network Models, Elsevier, 1995. [0-444-89292-3] 95 年 6 月購入

[Bar93] V. C. Barbosa, Massively Parallel Models of Computation, Ellis Horwood, 1993. [0-13-562968-3] 94 年 4 月購入

[BJK93] O. E. Barndorff-Nielsen, J. L. Jensen W. S. Kendall (eds.), Networks and Chaos – Statistical and Probabilistic Aspects, Chapman & Hall, 1993. [0-412-46530-2] 94 年 4 月購入

[BJS77] M. S. Bazaraa, J. J. Jarvis, H. D. Sherali, Linear Programming and Network Flows, Second Edition, John Wiley & Sons, 1977. [0-471-06015-1] 93 年 6 月購入

[BSS93] M. S. Bazaraa, H. D. Sherali, C. M. Shetty, Nonlinear Programming: Theory and Algorithms, Second Edition, John Wiley, 1993. [0-471-55793-5] 94 年 5 月購入

[Ber82] D. P. Bertsekas, Constrained Optimization and Lagrange Multiplier Methods, Academic Press, 1982. [0-12-093480-9] 93 年 6 月購入

[Ber87] D. P. Bertsekas, Dynamic Programming: Deterministic and Stochastic Models, Prentice-Hall, 1987. [0-13-221581-0] 95 年 2 月購入

[BT89] D. P. Bertsekas, J. N. Tsitsiklis, Parallel and Distributed Computation, Numerical Methods, Prentice-Hall, 1989. [0-13-648700-9] 93 年 9 月購入

[BESW94] J. Błazewicz, K. H. Ecker, G. Schmidt, J. Wȩglarz, Scheduling in Computer and Manufacturing Systems, Springer-Verlag, 1994. [3-540-58049-2] 95 年 1 月購入 [Ble90] G. E. Blelloch,

Vector Models for Data-Parallel Computing, The MIT Press, 1990. [0-262-02313-X] 93 年 8 月購入

[BS93] A. M. Borzyszkowski, S. Sokołowski (eds.), Mathematical Foundations of Computer science 1993, Springer-Verlag, 1993. [3-540-57182-5] 94 年 4 月購入

[BHM77] S. P. Bradley, A. C. Hax, T. L. Magnanti, Applied Mathematical Programming, Addison-Wesley, 1977. [0-201-00464-X] 93 年 7 月購入

[BK94] R. Bulirsch. D. Kraft (eds.), Computational Optimal Control, Birkhäuser Verlag, 1994. [3-7643-5105-6] 95 年 1 月購入

C [CU93] A. Cichocki, R. Unbehauen, Neural Networks for Optimization and Signal Processing, John Wiley, 1993. 93 年 6 月購入

[CL93] T. A. Ciriani, R. C. Leachman (eds.), Optimization Theory in Industry, John Wiley & Sons, 1993. [0-471-93492-5] 94 年 9 月購入

[CL94] T. A. Ciriani, R. C. Leachman (eds.), Optimization Theory in Industry 2, John Wiley & Sons, 1994. [0-471-94192-1] 94 年 9 月購入

[Cla90] F. H. Clarke, Optimization and Nonsmooth Analysis, SIAM, 1990. [0-89871-256-4] 94 年 9 月購入

[CLR92] E. G. Coffman, J., J. K. Lenstra, A. H. G. Rinnooy Kan, (eds.), Handbooks in Operations Research and Management Science, Vol. 3, Computing, North Holland, 1992. [0-444-88097-6] 93 年 7 月購入

[CL90] T. F. Coleman and Y. Li, (eds.), Large-Scale Numerical Optimization, SIAM, 1990. [0-89871-268-8] 94 年 2 月購入

[CPS92] R.W. Cottle, J.-S. Pang, R.E. Stone, The Linear Complementarity Problem, Academic Press, 1992. [0-12-192350-9] 93 年 12 月購入 [Cra91] R. E. Crandall,

Mathematica for the Sciences, Addison-Weslay, 1991. [0-201-51001-4] 94 年 5 月購入

D [Das95] M. S. Daskin, Network and Discrete Location: Models, Algorithms, and Applications, John Wiley & Sons, 1995. [0-471-01897-X] 95 年 10 月購入

[DSM94] Y. Davidor, H. P. Schwefel, R. Männer (eds.), Parallel Problem Solving from Nature – PPSN III, Springer-Verlag, 1994. [3-540-58484-6] 94 年 12 月購入

[Dav91] L. Davis (ed.), Handbook of Genetic Algorithms, Van Nostrand Reinhold, 1991. [0-442-00173-8] 94 年 6 月購入

[DS83] J. E. Dennis, Jr., R. B. Schnabel, Numerical Methods for Unconstrained Optimization and Nonlinear Equations, Prentice-Hall, 1983. [0-13-627216-9] 93 年 10 月購入

[DZ93] A. L. Dontchev, T. Zolezzi, Well-Posed Optimization Problems, Springer-Verlag, 1993. [3-540-56737-2] 94 年 2 月購入

[DP93] D. Z. Du, P. M. Pardalos (eds.), Network Optimization Problems: Algorithms, Applications and Complexity, World Scientific, 1993. [981-02-1277-1] 93 年 6 月購入

[DS94] D. Z. Du, J. Sun (eds.), Advances in Optimization and Approximation, Kluwer Academic Publishers, 1994. [0-7923-2785-3] 94 年 11 月購入

[DS94] D. Z. Du, J. Sun (eds.), Advances in Optimization and Approximation, Kluwer Academic Publishers, 1994. [0-7923-2785-3] 95 年 3 月購入 (2 冊目)

E [Els93] K.-H. Elster, Modern Mathematical Methods of Optimization, Akademie Verlag, 1993. [3-05-501452-9] 95 年 6 月購入 [EM92] J. R. Evans, E. Minieka,

Optimization Algorithms for Networks and Graphs: Second Edition, Revised and Expanded, Marcel Dekker, 1992. [0-8247-8062-5] 94 年 7 月購入

F [FM90] A. V. Fiacco, G. P. McCormick, Nonlinear Programming, SIAM, 1990. [0-89871-254-8] 94 年 3 月購入

[FF93a] A. E. Fincham, B. Ford (eds.), Parallel Computation, Oxford University Press, 1993. [0-19-853680-1] 94 年 4 月購入

[FF93b] A. E. Fincham, B. Ford (eds.), Parallel Computation, Oxford University Press, 1993. [0-19-853680-1] 94 年 7 月購入 (2 冊目)

[Fle87] R. Fletcher, Practical Methods of Optimization: second edition, John Wiley & Sons, 1987. [0-471-91547-5] 94 年 7 月購入

[FP92] C. A. Floudas, P. M. Pardalos (eds.), Recent Advances in Global Optimization, Princeton University Press, 1992. [0-691-08740-7] 94 年 6 月購入

[For93] S. Forrest (ed.), Proceedings of the fifth international conference on Genetic Algorithms, Morgan Kaufmann, 1993. [1-55860-299-2] 94 年 7 月購入

[FGK93] R. Fourer, D. M. Gay, B. W. Kernighan, AMPL: A Modeling Language for Mathematical Programming, The Scientific Press, 1993. [0-89426-232-7] 95 年 4 月購入

[Fou94] T. J. Fountain, Parallel computing: principles and practice, Cambridge University Press, 1994. [0-521-45131-0] 94 年 11 月購入

[FWM94] G. C. Fox, R. D. Wiliams, P. C. Messina (eds.), Parallel Computing Works!, Morgan Kaufmann, 1994. [1-55860-253-4] 94 年 7 月購入

G [GH93] W. Gander, J. Hřebíček, Solving Problems in scientific Computing Using MAPLE and MATLAB, Springer-Verlag, 1993. [3-540-57329-1] 94 年 8 月購入 [GJ79] M. R. Garey, D. S. Johnson,

Computers and Intractability: A Guide to the Theory of NP-Completeness, W. H. Freedman and Company, 1979. [0-7167-1045-5] 94 年 12 月購入

[GM95] F. Giannessi, A. Maugeri, Variational Inequalities and Network Equilibrium Problems, Plenum Press, 1995. [0-306-45007-0] 95 年 9 月購入

[GMW81] P. E. Gill, W. Murray, M. H. Wright, Practical Optimization, Academic Press, 1981. [0-12-283952-8] 93 年 6 月購入

[Gir94] C. Girault (ed.), Applications in Parallel and distributed Computing, North-Holland, 1994. [0-444-81870-7] 94 年 5 月購入

[Gla92] R. Glassey, Numerical Computing Using C, Academic Press, 1992. [0-12-286155-8] 93 年 6 月購入

[GMM94] P. A. Gloor, F. Makedon, J. W. Matthews, Parallel Computation: Practical Implementation of Algorithms and Machines, Springer-Verlag, 1994. [0-387-14213-4] 94 年 3 月購入

[GKP92] F. Glover, D. Klingman, N. V. Phillips, Network Models in Optimization and Their Applications in Practice, John Wiley, 1992. [0-471-57138-5] 93 年 6 月購入

[Gol89] D. E. Goldberg, Genetic Algorithms in Search, Optimization, and Machine Learning, Addison-Wesley, 1989. [0-201-15767-5] 94 年 9 月購入

[GGL94] G. Golub, A. Greenbaum, M. Luskin, Recent Advances in Iterative Methods, Springer-Verlag, 1994. [0-387-94252-1] 94 年 8 月購入

[GO93] G. H. Golub, J. M. Ortega, Scientific Computing: An Introduction with Parallel Computing, Academic Press, 1993. [0-12-289253-4] 93 年 6 月購入

[GV89] G. H. Golub, C. F. Van Loan, Matrix Computations : Second Edition, The Johns Hopkins University Press, 1989. [0-8018-3739-1]

93 年 8 月購入

[GG91] T. W. Gray, J. Glynn, Exploring Mathematics with Mathematica, Addison-Wesley, 1991. [0-201-52818-5] 94 年 5 月購入

[GG92] T. W. Gray, J. Glynn, The Beginner’s Guide to Mathematica : Version 2, Addison-Wesley, 1992. [0-201-58221-X] 94 年 5 月購入

[Gre93] H. J. Greenberg, A Computer-Assisted Analysis System for Mathematical Programming Models and Solutions : A User’s Guide for ANALYZE\begin{matrix}c \\ ◯ \end{matrix}, Kluwer Academic Press, 1993. [0-7923-9322-8] 94 年 1 月購入

[GL92] J. Gregory, C. Lin, Constrained Optimization in the Calculus of Variations and Optimal Control Theory, Van Nostrand Reinhold, 1992. [0-442-00722-1] 94 年 6 月購入

[GW94] D. F. Griffiths, G. A. Watson (eds.), Numerical Analysis 1993, Longman Scientific & Technical, 1994. [0-582-22569-X] 94 年 4 月購入

[GLS94] W. Gropp, E. Lusk, A. Skjellum, Using MPI: Portable Parallel Programming with the Message-Passing Interface, The MIT Press, 1994. [0-262-57104-8] 95 年 5 月購入

[GLS93] M. Grötshel, L. Lovász, A. Schrijver, Geometric Algorithms and Combinatorial Optimization: second corrected edition, Springer-Verlag, 1993. [3-540-56740-2] 94 年 7 月購入

H [HB93] S. Hariri, P. Bruce Berra, Proceedings of the 1993 International Conference on Parallel Processing: Vol. III Algorithms & Applications, CRC Press, 1993. [0-8493-8986-0] 94 年 4 月購入

[Har94] R. L. Harvey, Neural Network Principles, Prentice-Hall, 1994. [0-13-063330-5] 94 年 6 月購入 [HY94] J. Henry, J. P. Yvon (eds.),

System Modeling and Optimization, Springer-Verlag, 1994. [3-540-19893-8] 94 年 9 月購入

[HKP91] J. Hertz, A. Krogh, R. G. Palmer, Introduction to the theory of Neural Computation, Addison-Wesley, 1991. [0-201-51560-1] 94 年 6 月購入

[HS90] D. P. Heyman, M. J. Sobel (eds.), Stochastic Models, Elsevier Science Publishers, 1990. [0-444-87473-9] 94 年 10 月購入

[HL93a] J-B. Hiriart-Urruty, C. Lemaréchal, Convex analysis and Minimization Algorithms I: Fundamentals, Springer-Verlag, 1993. [3-540-56850-6] 94 年 2 月購入

[HL93b] J-B. Hiriart-Urruty, C. Lemaréchal, Convex analysis and Minimization Algorithms II: Advanced Theory and Bundle Methods, Springer-Verlag, 1993. [3-540-56852-2] 94 年 2 月購入

[Hol92] J. H. Holland, Adaptation in Natural and Artificial Systems, The MIT Press, 1992. [0-262-58111-6] 95 年 1 月購入

[HP95] R. Horst, P. M. Pardalos (eds.), Handbook of Global Optimization, Kluwer Academic Publishers, 1995. [0-7923-3120-6] 95 年 3 月購入

[HPT95] R. Horst, P. M. Pardalos, N. V. Thoai, Inroduction to Global Optimization, Kluwer Academic Publishers, 1995. [0-7923-3557-0] 95 年 8 月購入

I [IC94] J. P. Ignizio, T. M. Cavalier, Linear Programming, Prentice-Hall, 1994. [0-13-183757-5] 94 年 7 月購入

[Isa92] G. Isac, Complementarity Problems, Springer-Verlag, 1992. [3-540-56251-6] 93 年 12 月購入

J [Jah94] J. Jahn, Introduction to the Theory of Nonlinear Optimization, Springer-Verlag, 1994. [3-540-58320-3] 94 年 11 月購入 K [KW94] P. Kall, S. W. Wallace,

Stochastic Programming, John Wiley & Sons, 1994. [0-471-95158-7] 94 年 9 月購入

[KA93] A. Kłos, P. W. Aitchison, Network Solution: Theory and Practice, Ellis Horwood, 1993. [0-13-613589-7] 94 年 3 月購入

[Knu84] D. E. Knuth, The TEXbook, Addison-Wesley, 1984. [0-201-13448-9] 94 年 9 月購入

[KRS94] S. Komlósi, T. Rapcsák, S. Schaible (eds.), Generalized Convexity, Springer, 1994. [3-540-57624-X] 94 年 7 月購入

[KT93] A. P. Korostelev, A. B. Tsybakov, Minimax Theory of Image Reconstruction, Springer, 1993. [3-540-94028-6] 93 年 6 月購入

[KG93] J. S. Kowalik, L. Grandinetti (eds.), Software for Parallel Computation, Springer, 1993. [3-540-56451-9] 93 年 7 月購入

[KGK94] R. Kruse, J. Gebhardt, R. Klawonn, Foundations of Fuzzy Systems, John Wiley & Sons, 1994. [0-471-94243-X] 94 年 11 月購入

[KGGK94] V. Kumar, A. Grama, A. Gupta, G. Karypis, Introduction to Parallel Computing: Design and Analysis of Algorithms, The Benjamin/Cummings Publishing Company, 1994. [0-8053-3170-0]

94 年 6 月購入

L [Lam86] L. Lamport, LaTEX A Document Preparation System, Addison-Wesley, 1986. [0-201-15790-X] 94 年 9 月購入

[Lan89] C. G. Langton (ed.), Artificial Life, Addison-Wesley, 1989. [0-201-09356-1] 94 年 9 月購入

[LTFR92] C. G. Langton, C. Talor, J. D. Farmer, S. Rasmussen (eds.), Artificial Life II, Addison-Wesley, 1992. [0-201-52570-4] 94 年 9 月購入 [Lan94] C. G. Langton (ed.),

Artificial Life III, Addison-Wesley, 1994. [0-201-62492-3] 94 年 7 月購入

[Law76] E. L. Lawler, Combinatorial Optimization: Networks and Matroids, Sounders College Publishing, 1976. [0-03-084866-0] 94 年 6 月購入

[LLRS85] E. L. Lawler, J. K. Lenstra, A. H. G. Rinnooy Kan, D. B. Shmoys (eds.), The Traveling Salesman Problem: A Guided Tour of Combinatorial Optimization, John Wiley & Sons, 1985. [0-471-90413-9] 94 年 7 月購入

[Lev94] E. S. Levitin, Perturbation Theory in Mathematical Programming and its Applications, John Wiley & Sons, 1994. [0-471-93935-8] 94 年 9 月購入

[LSP93] Th. Lippert, K. Schilling, P. Ueberholz, Science on the Connection Machine, World Scientific, 1993. [981-02-1206-2] 93 年 7 月購入

[Lue84] D. G. Luenberger, Linear and Nonlinear Programming: Second Edition, Addison-Wesley, 1984. [0-201-15794-2] 94 年 9 月購入

M [MRP86] J. L. McClelland, D. E. Rumelhaut and the PDP Research Group, Parallel Distributed Processing Volume 2: Psychological and Biological Methods, The MIT Press, 1986. [0-262-63110-5] 94 年 2 月購入

[Mes91] J. P. Mesirov (ed.), Very Large Scale Computation in the 21st Century, SIAM, 1991. [0-89871-279-3] 93 年 9 月購入

[Mic94] Z. Michalewicz, Genetic algorithms + Data Structures = Evolution Programs (Second, Extended Edition), Springer-Verlag, 1994. [3-540-58090-5] 94 年 11 月購入

[MP88] M. L. Minsky and S. A. Papert, Perceptrons (Expanded Edition), The MIT Press, 1988. [0-262-63111-3] 93 年 8 月購入 [MF90] P. B. Mirchandani, R. L. Francis (eds.),

Discrete Location Theory, John Wiley & Sons, 1990. [0-471-89233-5] 95 年 10 月購入

[MW93] J. J. Moré and S. J. Wright, Optimization Software Guide, SIAM, 1993. [0-89871-322-6] 94 年 2 月購入

[MP93] T. E. Morton, D. W. Pentico, Heuristic Scheduling Systems, John Wiley & Sons, 1993. [0-471-57819-3] 95 年 1 月購入

N [Nag93] A. Nagurney, Network Economics: A Variational Inequality Approach, Advances in Computational Economics 1, Kluwer Academic Publishers, 1993. [0-7923-9293-0] 93 年 12 月購入

[Naz94] J. L. Nazareth, The Newton-Cauchy Framework, Springer-Verlag, 1994. [3-540-57671-1] 94 年 8 月購入

[NRT89] G. L. Nemhauser, A. H. G. Rinnooy Kan, M. J. Todd, Optimization, North-Holland, 1989. [0-444-87284-1] 94 年 5 月購入

[NW88] G. L. Nemhauser, L. A. Wolsey, Integer and Combinatorial Optimization, John Wiley & Sons, 1988. [0-471-82819-X] 94 年 7 月購入

[NT93] E. D. Nering, A. W. Tucker, Linear Programs and Related Problems, Academic Press, 1993. [0-12-515440-2] 93 年 6 月購入

[NN94a] Y. Nesterov, A. Nemirovskii, Interior-Point Polynomial Algorithms in Convex Programming, SIAM, 1994. [0-89871-319-6] 94 年 3 月購入

[NN94b] Y. Nesterov, A. Nemirovskii, Interior-Point Polynomial Algorithms in Convex Programming, SIAM, 1994. [0-89871-319-6] 94 年 6 月購入 (2 冊目)

O [OR70] J.M. Ortega, W.C. Rheinboldt, Iterative Solution of Nonlinear Equations in Several Variables, Academic Press, 1970. [0-12-528550-7] 94 年 1 月購入 P [Pad95] M. Padberg,

Linear Optimization and Extensions, Springer-Verlag, 1995. [3-540-58734-9] 95 年 5 月購入

[PGW93] G. F. Page, J. B. Gomm, D. Williams (eds.), Application of Neural Networks to Modeling and Control, Chapman & Hall, 1993. [0-412-54760-0] 94 年 4 月購入

[Pap94] C. H. Papadimitriou, Computational Complexity, Addison-Wesley, 1994. [0-201-53082-1] 94 年 4 月購入

[PS82] C. H. Papadimitriou, K. Steiglitz, Combinatorial Optimization: Algorithms and Complexity, Prentice-Hall, 1982. [0-13-152462-3] 93 年 9 月購入

[Par93] P. M. Pardalos (ed.), Complexity in Numerical Optimization, World Scientific, 1993. [981-02-1415-4] 93 年 10 月購入

[PW94] P. M. Pardalos, H. Wolkowicz (eds.), Quadratic Assignment and Related Problems, American Mathematical Society, 1994. [0-8218-6607-9] 94 年 12 月購入

[PSA95] D. W. Pearson, N. C. Steele, R. F. Albrecht (eds.), Artificial Neural Nets and Genetic Algorithms, Springer-Verlag, 1995. [3-211-82692-0] 95 年 9 月購入

[PJS92] H. O. Peitgen, H. Jürgens, D. Saupe, Chaos and Fractals: New Frontiers of Science, Springer-Verlag, 1992. 93 年 6 月購入

[Pit93] I. Pitas (ed.), Parallel Algorithms for Digital Image Processing, Computer Vision and Neural Networks, John Wiley, 1993. 93 年 6 月購入

[Pol87] B. T. Polyak, Introduction to Optimization, Optimization Software, 1987. [0-911575-14-6] 94 年 9 月購入

[PRV94] I. Prívara, B. Rovan, P. Ružička (eds.), Mathematical Foundations of Computer Science 1994, Springer-Verlag, 1994. [3-540-58338-6] 94 年 11 月購入 [Psh94] B. N. Pshenichnyj,

The Linearization Method for Constrained Optimization, Springer-Verlag, 1994. [3-540-57037-3] 94 年 7 月購入

[Put94] M. L. Puterman, Markov Decision Processes: Discrete Stochastic Dynamic Programming, John Wiley & Sons, 1994. [0-471-61977-9] 94 年 10 月購入

R [Raw91] G. J. E. Rawlins (ed.) Foundations of Genetic Algorithms, Morgan Kaufmann Publishers, 1991. [1-55860-170-8] 93 年 7 月購入

[Rei94] G. Reinelt, The Traveling Salesman: Computational Solutions for TSP Applications, Springer-Verlag, 1994. [3-540-58334-3] 94 年 9 月購入

[Ric93] J. R. Rice, Numerical Methods, Software, and Analysis: Second Edition, Academic Press, 1993. [0-12-587755-2] 93 年?月購入

[Roc70] R. T. Rockafellar, Convex Analysis Princeton University Press, 1970. [0-691-08069-0] 93 年 9 月購入

[Roc74] R. T. Rockafellar, Conjugate Duality and Optimization, SIAM, 1974. [0-89871-013-8] 94 年 12 月購入

[Roc84] R. T. Rockafellar, Network Flows and Monotropic Optimization, John Wiley & Sons, 1984. [0-471-88078-7] 94 年 8 月購入

[Roc84] R. T. Rockafellar, Network Flows and Monotropic Optimization, John Wiley & Sons, 1984. [0-471-88078-7] 94 年 12 月購入 (2 冊目)

[RMP86] D. E. Rumelhaut, J. L. McClelland and the PDP Research Group, Parallel Distributed Processing Volume 1: Foundations, The MIT Press, 1986. [0-262-68053-X] 94 年 2 月購入

S [Spe94] H.-P. Schwefel, Evolution and Optimum Seeking, John Wiley & Sons, 1995. [0-471-57148-2] 95 年 6 月購入 [Spe94] E. Spedicato (ed.),

Algorithms for Continuous Optimization: The State of the Art, Kluwer Academic Publishers, 1994. [0-7923-2859-0] 94 年 6 月購入

T [TGS94] V. S. Tanaev, V. S. Gordon, Y. M. Shafransky, Scheduling Theory. Single-Stage Systems, Kluwer Academic Publishers, 1994. [0-7923-2853-1] 95 年 2 月購入

[TSS94] V. S. Tanaev, Y. N. Sotskov, V. A. Strusevich, Scheduling Theory. Multi-Stage Systems, Kluwer Academic Publishers, 1994. [0-7923-2854-X] 95 年 2 月購入

[Tar83] R. E. Tarjan, Data Structures and Network Algorithms, SIAM, 1983. [0-89871-187-8] 94 年 6 月購入

[Tij94] H. C. Tijms, Stochastic Models: An Algorithmic Approach, John Wiley & Sons, 1994. [0-471-94380-0] 95 年 1 月購入

[TIM93] THE TIMES ATLAS OF THE WORLD, TIMES BOOKS, 1993. [0-7230-0492-7] 94 年 5 月購入

V [VW92] J. H. van Lint, R. M. Wilson, A Course in Combinatorics, Cambridge University Press, 1992. [0-521-42260-4] 93 年 6 月購入

W [Whi93] L. D. Whitley (ed.), Foundations of Genetic Algorithms · 2, Morgan Kaufmann Publishers, 1993. [1-55860-263-1] 93 年 6 月購入

[Wil90] H. P. Williams, Model Building in Mathematical Programming: Third edition Revised, John Wiley & Sons, 1990. [0-471-94111-5] 94 年 10 月購入

Z [Zhi91] A. A. Zhiglijavsky, Theory of Global Random Search, Kluwer Academic Publishers, 1991. [0-7923-1122-1] 94 年 8 月購入